If P (A|B) = 0.5, P (A) = 0.45 and P (B) = ?


1
Expert's answer
2022-05-10T08:30:51-0400

Bayes' theorem:

"P(A|B)=\frac{P(B|A)P(A)}{P(B)}"

then:

"P(B)=\frac{P(B|A)P(A)}{P(A|B)}=\frac{0.45P(B|A)}{0.5}=0.9P(B|A)"